# Towards autonomous photogrammetric forest inventory using a lightweight under-canopy robotic drone

### v3

@@ 3.1 Autonomous navigation algorithms
Among the open source solutions presented in Section 2.2, the solution by Zhou et al. (2022) was chosen as the basis of the prototype. The solution was the only open source camera-based solution that was proven to work with only one stereo camera, making it possible to build a smaller and lighter platform. Furthermore, the test environment reported in the original article (dense bamboo forest) was the most difficult among the proposed solutions, although it had a highly different structure from typical boreal forests.
@@
Figure 1: The system architecture used in the data collection. In comparison to the original EGO-Planner-v2 article by Zhou et al. (2022), UWB-based Drift correction and Wireless trajectory broadcast network modules used for swarm navigation were omitted in this study. Stereo images were downsampled for VIO to guarantee real-time estimation. Full-resolution stereo images were also saved and post-processed to point clouds used in the DBH estimation.
@@ 3.1.1 EGO-Planner-v2
The source code of the planner algorithm by Zhou et al. (2022) was named on GitHub as EGO-Planner-v2 (ZJU FAST Lab Team 2022a). Although EGO-Planner-v2 is designed for swarm flying, in this study, the algorithm was modified for a single drone system to focus on navigation performance in challenging forest environments. The planner runs on top of the Robot Operating System (ROS) (Quigley et al. 2009). The mapping of EGO-Planner-v2 is based on probabilistic occupancy grid maps (Moravec and Elfes 1985). The mapping module uses depth images as input and the system adopts the fixed-size circular buffers proposed by Usenko et al. (2017) to maintain the local map. The maps also have a virtual floor and a virtual ceiling defining the minimum and maximum altitudes of the grid map, respectively. The virtual floor and the virtual ceiling restrict the altitude at which the algorithm can plan the paths.
In trajectory planning, EGO-Planner-v2 uses a MINCO (minimum control) (Wang et al. 2022) trajectory representation designed for differentially flat systems, such as quadrotors (Mellinger and Kumar 2011). The planned trajectories are represented as piece-wise polynomial splines with decoupled temporal and spatial parameters (Zhou et al. 2022). The trajectories are optimized iteratively by minimizing the weighted sum of metrics defined for trajectory smoothness and flying time. The feasibility of the planned trajectory is forced by deforming the shape to avoid obstacles and by restricting the magnitudes of the trajectory velocity, acceleration, and jerk to meet the dynamical constraints of the drone. For computational efficiency, trajectory constraints are enforced through integrals of penalty functions with large penalty weights, and the integrals are evaluated by a finite sum of equally spaced samples along the timeline. In a simplified form, the optimization of time-dependent objectives $J$ in one piece of the trajectory can be represented by
@@
The trajectory planning is continuously performed within a defined local planning distance towards a user-defined global goal until the goal is reached. Since the safety constraints are modeled as integrals in the penalty functions, the feasibility of the trajectory needs to be verified with a post-check after planning (Zhou et al. 2022). Due to the non-static probabilistic map, the post-check after trajectory planning guarantees feasibility only at a certain moment. Hence, a collision-check process is continuously running in the background. If that process detects a potential collision, a trajectory replanning is activated immediately. If the replanning fails and the predicted time to collision is under a given threshold, the planner performs an emergency stop. After the drone has stopped, the trajectory planner tries to start up again and find a new feasible trajectory.
@@ 3.1.2 VINS-Fusion
The Visual-inertial odometry (VIO) algorithm used by Zhou et al. (2022) was VINS-Fusion (Qin, Li, and Shen 2018; Qin et al. 2019b)(VINS stands for ’Visual-Inertial system’), which used grayscale stereo images and IMU data to track the position and orientation of the drone.
VINS-Fusion detects Shi-Tomasi corner features (Shi and Tomasi 1994), and the features are tracked between frames with Kanade–Lucas–Tomasi (KLT) Tracker (Lucas and Kanade 1981). IMU data is used for preintegration, i.e., position, velocity, and orientation are integrated from IMU measurements between two frames. The state optimization is performed within a sliding window, including the position, velocity, rotation, and IMU biases. To reduce the computational complexity, old measurements are marginalized into a prior of the estimation process. Certain frames are also selected as keyframes, which are saved to a global pose graph after being marginalized out from the sliding window. Keyframes are selected by considering the number of tracked features and the average parallax between these features in the most recent camera frame and the latest keyframe.
VINS-Fusion supports loop detection-based relocalization and global pose graph optimization to reduce the accumulated drift in the pose estimate. The loops are detected by storing additional corner features as BRIEF descriptors (Calonder et al. 2010) to a DBoW2 (Galvez-López and Tardos 2012) bag-of-word database. After a loop detection, the relocalization process aligns the latest window to past poses and optimizes it utilizing also the feature correspondences with past poses. In addition to sliding-window relocalization, the global pose graph is also optimized after loop detections to enforce the global consistency of the poses. Since the roll and pitch angle estimates are drift accumulation-free due to full observability from the gravity measurement of the IMU, the pose graph optimization is performed in 4 degrees of freedom.
VINS-Fusion also adopts an online temporal calibration method for inertial and visual measurements (Qin and Shen 2018). The method can be used to adjust the time offset $t_{d}$ between the camera and the IMU online.
In the swarm operations of the original article (Zhou et al. 2022), the drones used also relative distances measured with Ultra Wide Band (UWB) sensors to correct the drift in the pose estimate of VINS-Fusion. However, with a single drone system in this study, that function is omitted.
@@ 3.1.3 Tracking controller
The original EGO-Planner-v2 article (Zhou et al. 2022) did not specify the method for trajectory tracking. In this study, an open source tracking controller (ZJU FAST Lab Team 2022b) by Zhejiang University was used. The tracking controller gets setpoints for position, velocity, acceleration, and yaw from the trajectory planner and converts those to attitude and thrust setpoints. The setpoints are forwarded to the PX4 flight controller software (PX4 Community 2023) responsible for the low-level attitude and angular rate controls.
@@ 3.2 Drone hardware
In the physical robot drone prototype, an onboard computer NVIDIA Jetson Orin NX (Santa Clara, California, USA) was used. The used PX4-compliant autopilot was Holybro Kakute H7 (Hong Kong, China), which also contained the IMU. The navigation algorithms mentioned in the previous subsection ran on top of the ROS, and the MAVROS protocol was used for communication between the onboard computer and the autopilot. As a camera, Intel RealSense D435 (Intel Realsense Development Team 2023)(Keselman et al. 2017) was used. The camera was capturing both depth images for 3D mapping and grayscale stereo images for VIO.
@@
Stereoscopic RealSense D435 has an infrared laser projector emitting dots, improving the triangulation in low light conditions and for homogeneous surfaces. However, the dots of the laser emitter cause problems for VIO since they are moving with the camera. Therefore, in this study, the infrared emitter was disabled. Since sunlight contains infrared light, which interferes with the artificial features emitted (Kuan, Ee, and Wei 2019), and forests typically do not contain homogeneous surfaces, disabling the emitter should not affect the camera performance in outdoor conditions.
The camera-IMU system was calibrated with an open source camera calibration toolbox, kalibr (Furgale, Barfoot, and Sibley 2012; Furgale, Rehder, and Siegwart 2013; Maye, Furgale, and Siegwart 2013). In the calibration, the intrinsic and extrinsic parameters of the left and right grayscale cameras and the time shift between the camera and the autopilot IMU were estimated from a recorded data sequence, where Aprilgrid (Olson 2011) was used as a calibration target. The IMU noise parameters needed for the calibration were first obtained with the open source tool Allan Variance ROS (Oxford Dynamic Robot Systems Group 2022), and the values obtained were multiplied by 10 before the calibration as suggested in the wiki of the kalibr project (ETH Zürich Autonomous Systems Lab 2023).
@@ 3.3 Test sites
The performance of the system was evaluated in three different boreal forest areas. The flight tests of the system were performed in two test areas located in Evo, Finland (61.19°N, 25.11°E). The flight paths went through well-documented forest sample plots that have been widely used in previous studies, such as in Liang et al. (2018, 2019); Wang et al. (2019); Hyyppä et al. (2020a, c), and are classified into three complexity categories, i.e., ’easy’, ’medium’, and ’difficult’ based on, mainly tree density and amount of vegetation under the canopy in the plots. The third test area, a spruce forest in Paloheinä, Finland (60.26°N, 24.92°E) was used only to test the loop detection performance of VINS-Fusion.
@@
For reference purposes, circular GCPs provided by Agisoft Metashape (Agisoft LLC, St. Petersburg, Russia) were placed within the forest plots used in the flight tests. The targets had a diameter of 24.7 cm and the middle dots had a diameter of 7 cm. Relative distances between the GCPs were measured with a Leica Nova TS60 total station with a precision of approximately 2 mm (Leica Geosystems 2020). Figure 5 shows the GCPs in ’Evo-difficult’ during the test flights.
@@ 3.5 Structure-from-Motion Multi-View Stereo based reconstruction
Stereo image processing was carried out using Agisoft Metashape software (Agisoft LLC, St. Petersburg, Russia) (Metashape Development Team 2021) following similar processing steps as introduced by Viljanen et al. (2018). Every sixth stereo pair was used in the Metashape processing with a stereo baseline of 5 cm and an accuracy setting of ±0.0015 m. In the image orientation processing, the quality setting was set to high, meaning that two times downsampled images were used. The settings for the number of key points and tie points per image were 40,000 and 4,000, respectively. To enhance the quality of the post-processed reference trajectories, GCPs were used in the processing with an accuracy setting of $\pm 0.005$ m. Automatic outlier removal was performed using the gradual selection tools of the software based on re-projection error, reconstruction uncertainty, and projection accuracy. Subsequently, a bundle adjustment was carried out, involving optimization of the sparse point cloud along with the interior (IOP) and the exterior orientation parameters (EOP). Following that, a dense point cloud was calculated using the highest quality setting and mild depth filtering. Finally, the EOPs and dense point clouds were exported.
@@ 3.6 Tree parameter estimation
In this study, trees were detected, and their DBHs were estimated using an automatic stem diameter estimation algorithm proposed by Hyyppä et al. (2020b) and further developed by Hyyppä et al. (2020a) to be suitable for a larger variety of point clouds, including photogrammetric point clouds, and point clouds produced with 3D mobile laser scanners and TLS. The method described by Hyyppä et al. (2020a) was applied to estimate the DBH from point clouds collected with the drone and a TLS system, the latter of which provided reference measurements of DBH (see details in Section 3.7).
The point cloud was first segmented into 40-cm-tall (20-cm-tall for TLS) vertical segments and 5-s-long temporal segments. The temporal segmentation was done to mitigate point cloud distortions caused by potential slowly-accumulating positioning errors of the moving drone system, and thus, the temporal segmentation was not performed for the TLS data. The points potentially corresponding to the tree stems were detected by clustering the points with density-based clustering (DBSCAN, (Ester et al. 1996)) and filtered by fitting a circle to each cluster while accepting only those clusters satisfying the following heuristic criteria: the cluster contained more than 35 points and at least 80 % of the points within 30 mm of the circular arc, the fitted circle radius was between 4 cm and 40 cm, and the points covered at least a 60° central angle of the fitted circle. The accepted clusters were further grouped into tree stems by applying DBSCAN for the center points of the clusters, and the growth direction of each tree was estimated with principal component analysis (PCA). The points within accepted clusters were projected into the plane perpendicular to the growth direction for the final diameter estimation. The diameters were processed into a stem curve by automatically removing clear outliers, followed by fitting a smoothing spline to the remaining diameter estimates as a function of height. Finally, the DBH estimates were calculated by interpolating the stem curve at a height of 1.3 m above the ground. For a detailed description of the method, the reader is referred to Hyyppä et al. (2020b, a).
@@ 3.7.1 Autonomous flying performance
The accuracy of the VIO localization was evaluated by calculating the Absolute Trajectory Error (ATE), as proposed by Sturm et al. (2012), with respect to the post-processed SfM reference trajectories. In this study, ATE was calculated only for the position estimate, and the error in the orientation estimate was not evaluated. ATE was calculated using the trajectory evaluation toolbox by Zhang and Scaramuzza (2018), which aligns the trajectory estimates with the reference trajectories with the method by Umeyama (1991). After aligning the trajectories, the toolbox calculates the RMSE of the individual state errors to generate a single metric error value representing the accuracy of the entire trajectory. In the mathematical form, the error can be presented as
@@
In the first tests conducted in a simulator as described by (Karjalainen et al. 2023), the VINS-Fusion estimate was found to be more accurate when the $t_{d}$ value was fixed to a constant, compared to when the online estimation method (Qin and Shen 2018) was used. For this reason, during the test flights, the $t_{d}$ value was fixed to the one obtained from kalibr. However, the calibration estimate is never perfect, and the prototype did not have actual hardware synchronization between the camera and the autopilot IMU. For that reason, after the test flights, the VINS-Fusion estimation was also run with online estimated $t_{d}$ with the camera and IMU data recorded during the flight tests. The estimates were performed with the onboard computer of the drone, and the $t_{d}$ estimate obtained with kalibr was used as the initial value of the online estimation.

### v4

@@ 3.1 Autonomous navigation algorithms
Among the open source solutions presented in Section 2.2, the solution by zhou2022swarm was chosen as the basis of the prototype. The solution was the only open source camera-based solution that was proven to work with only one stereo camera, making it possible to build a smaller and lighter platform. Furthermore, the test environment reported in the original article (dense bamboo forest) was the most difficult among the proposed solutions, although it had a highly different structure from typical boreal forests.
@@
Figure 1: The system architecture used in the data collection. In comparison to the original EGO-Planner-v2 article by zhou2022swarm, UWB-based Drift correction and Wireless trajectory broadcast network modules used for swarm navigation were omitted in this study. Stereo images were downsampled for VIO to guarantee real-time estimation. Full-resolution stereo images were also saved and post-processed to point clouds used in the DBH estimation.
@@ 3.1.1 EGO-Planner-v2
The source code of the planner algorithm by zhou2022swarm was named on GitHub as EGO-Planner-v2 (EGO-Planner-code). Although EGO-Planner-v2 is designed for swarm flying, in this study, the algorithm was modified for a single drone system to focus on navigation performance in challenging forest environments. The planner runs on top of the Robot Operating System (ROS) (quigley2009ros). The mapping of EGO-Planner-v2 is based on probabilistic occupancy grid maps (moravec1985high). The mapping module uses depth images as input and the system adopts the fixed-size circular buffers proposed by usenko2017 to maintain the local map. The maps also have a virtual floor and a virtual ceiling defining the minimum and maximum altitudes of the grid map, respectively. The virtual floor and the virtual ceiling restrict the altitude at which the algorithm can plan the paths.
In trajectory planning, EGO-Planner-v2 uses a MINCO (minimum control) (MINCO) trajectory representation designed for differentially flat systems, such as quadrotors (differentialFlatness). The planned trajectories are represented as piece-wise polynomial splines with decoupled temporal and spatial parameters (zhou2022swarm). The trajectories are optimized iteratively by minimizing the weighted sum of metrics defined for trajectory smoothness and flying time. The feasibility of the planned trajectory is forced by deforming the shape to avoid obstacles and by restricting the magnitudes of the trajectory velocity, acceleration, and jerk to meet the dynamical constraints of the drone. For computational efficiency, trajectory constraints are enforced through integrals of penalty functions with large penalty weights, and the integrals are evaluated by a finite sum of equally spaced samples along the timeline. In a simplified form, the optimization of time-dependent objectives $J$ in one piece of the trajectory can be represented by
@@
The trajectory planning is continuously performed within a defined local planning distance towards a user-defined global goal until the goal is reached. Since the safety constraints are modeled as integrals in the penalty functions, the feasibility of the trajectory needs to be verified with a post-check after planning (zhou2022swarm). Due to the non-static probabilistic map, the post-check after trajectory planning guarantees feasibility only at a certain moment. Hence, a collision-check process is continuously running in the background. If that process detects a potential collision, a trajectory replanning is activated immediately. If the replanning fails and the predicted time to collision is under a given threshold, the planner performs an emergency stop. After the drone has stopped, the trajectory planner tries to start up again and find a new feasible trajectory.
@@ 3.1.2 VINS-Fusion
The Visual-inertial odometry (VIO) algorithm used by zhou2022swarm was VINS-Fusion (qin2018vins; qin2019general)(VINS stands for ’Visual-Inertial system’), which used grayscale stereo images and IMU data to track the position and orientation of the drone.
VINS-Fusion detects Shi-Tomasi corner features (shiTomasi1994), and the features are tracked between frames with Kanade–Lucas–Tomasi (KLT) Tracker (lucas1981iterative). IMU data is used for preintegration, i.e., position, velocity, and orientation are integrated from IMU measurements between two frames. The state optimization is performed within a sliding window, including the position, velocity, rotation, and IMU biases. To reduce the computational complexity, old measurements are marginalized into a prior of the estimation process. Certain frames are also selected as keyframes, which are saved to a global pose graph after being marginalized out from the sliding window. Keyframes are selected by considering the number of tracked features and the average parallax between these features in the most recent camera frame and the latest keyframe.
VINS-Fusion supports loop detection-based relocalization and global pose graph optimization to reduce the accumulated drift in the pose estimate. The loops are detected by storing additional corner features as BRIEF descriptors (BRIEF) to a DBoW2 (DBoW2) bag-of-word database. After a loop detection, the relocalization process aligns the latest window to past poses and optimizes it utilizing also the feature correspondences with past poses. In addition to sliding-window relocalization, the global pose graph is also optimized after loop detections to enforce the global consistency of the poses. Since the roll and pitch angle estimates are drift accumulation-free due to full observability from the gravity measurement of the IMU, the pose graph optimization is performed in 4 degrees of freedom.
VINS-Fusion also adopts an online temporal calibration method for inertial and visual measurements (onlineTd). The method can be used to adjust the time offset $t_{d}$ between the camera and the IMU online.
In the swarm operations of the original article (zhou2022swarm), the drones used also relative distances measured with Ultra Wide Band (UWB) sensors to correct the drift in the pose estimate of VINS-Fusion. However, with a single drone system in this study, that function is omitted.
@@ 3.1.3 Tracking controller
The original EGO-Planner-v2 article (zhou2022swarm) did not specify the method for trajectory tracking. In this study, an open source tracking controller (px4ctrl) by Zhejiang University was used. The tracking controller gets setpoints for position, velocity, acceleration, and yaw from the trajectory planner and converts those to attitude and thrust setpoints. The setpoints are forwarded to the PX4 flight controller software (PX4code) responsible for the low-level attitude and angular rate controls.
@@ 3.2 Drone hardware
In the physical robot drone prototype, an onboard computer NVIDIA Jetson Orin NX (Santa Clara, California, USA) was used. The used PX4-compliant autopilot was Holybro Kakute H7 (Hong Kong, China), which also contained the IMU. The navigation algorithms mentioned in the previous subsection ran on top of the ROS, and the MAVROS protocol was used for communication between the onboard computer and the autopilot. As a camera, Intel RealSense D435 (realsense)(realsense_article) was used. The camera was capturing both depth images for 3D mapping and grayscale stereo images for VIO.
@@
Stereoscopic RealSense D435 has an infrared laser projector emitting dots, improving the triangulation in low light conditions and for homogeneous surfaces. However, the dots of the laser emitter cause problems for VIO since they are moving with the camera. Therefore, in this study, the infrared emitter was disabled. Since sunlight contains infrared light, which interferes with the artificial features emitted (IRoutdoor), and forests typically do not contain homogeneous surfaces, disabling the emitter should not affect the camera performance in outdoor conditions.
The camera-IMU system was calibrated with an open source camera calibration toolbox, kalibr (kalibr3; kalibr2; kalibr1). In the calibration, the intrinsic and extrinsic parameters of the left and right grayscale cameras and the time shift between the camera and the autopilot IMU were estimated from a recorded data sequence, where Aprilgrid (apriltag) was used as a calibration target. The IMU noise parameters needed for the calibration were first obtained with the open source tool Allan Variance ROS (AllanVarianceROS), and the values obtained were multiplied by 10 before the calibration as suggested in the wiki of the kalibr project (kalibrWikiIMU).
@@ 3.3 Test sites
The performance of the system was evaluated in three different boreal forest areas. The flight tests of the system were performed in two test areas located in Evo, Finland (61.19°N, 25.11°E). The flight paths went through well-documented forest sample plots that have been widely used in previous studies, such as in liang2018situ; liang2019forest; wang2019field; hyyppa2020under; hyyppa2020comparison, and are classified into three complexity categories, i.e., ’easy’, ’medium’, and ’difficult’ based on, mainly tree density and amount of vegetation under the canopy in the plots. The third test area, a spruce forest in Paloheinä, Finland (60.26°N, 24.92°E) was used only to test the loop detection performance of VINS-Fusion.
@@
For reference purposes, circular Ground Control Points (GCPs) provided by Agisoft Metashape (Agisoft LLC, St. Petersburg, Russia) were placed within the forest plots used in the flight tests. The targets had a diameter of 24.7 cm and the middle dots had a diameter of 7 cm. Relative distances between the GCPs were measured with a Leica Nova TS60 total station with a precision of approximately 2 mm (TS60Datasheet). Figure 5 shows the GCPs in ’Evo-difficult’ during the test flights.
@@ 3.5 Structure-from-Motion Multi-View Stereo based reconstruction
Stereo image processing was carried out using Agisoft Metashape software (Agisoft LLC, St. Petersburg, Russia) (Metashape) following similar processing steps as introduced by Viljanen2018. Every sixth stereo pair was used in the Metashape processing with a stereo baseline of 5 cm and an accuracy setting of ±0.0015 m. In the image orientation processing, the quality setting was set to high, meaning that two times downsampled images were used. The settings for the number of key points and tie points per image were 40,000 and 4,000, respectively. To enhance the quality of the post-processed reference trajectories, GCPs were used in the processing with an accuracy setting of $\pm 0.005$ m. Automatic outlier removal was performed using the gradual selection tools of the software based on re-projection error, reconstruction uncertainty, and projection accuracy. Subsequently, a bundle adjustment was carried out, involving optimization of the sparse point cloud along with the interior (IOP) and the exterior orientation parameters (EOP). Following that, a dense point cloud was calculated using the highest quality setting and mild depth filtering. Finally, the EOPs and dense point clouds were exported.
@@ 3.6 Tree parameter estimation
In this study, trees were detected, and their DBHs were estimated using an automatic stem diameter estimation algorithm proposed by hyyppa2020accurate and further developed by hyyppa2020under to be suitable for a larger variety of point clouds, including photogrammetric point clouds, and point clouds produced with 3D mobile laser scanners and TLS. The method described by hyyppa2020under was applied to estimate the DBH from point clouds collected with the drone and a TLS system, the latter of which provided reference measurements of DBH (see details in Section 3.7).
The point cloud was first segmented into 40-cm-tall (20-cm-tall for TLS) vertical segments and 5-s-long temporal segments. The temporal segmentation was done to mitigate point cloud distortions caused by potential slowly-accumulating positioning errors of the moving drone system, and thus, the temporal segmentation was not performed for the TLS data. The points potentially corresponding to the tree stems were detected by clustering the points with density-based clustering (DBSCAN, (ester1996DBSCAN)) and filtered by fitting a circle to each cluster while accepting only those clusters satisfying the following heuristic criteria: the cluster contained more than 35 points and at least 80 % of the points within 30 mm of the circular arc, the fitted circle radius was between 4 cm and 40 cm, and the points covered at least a 60° central angle of the fitted circle. The accepted clusters were further grouped into tree stems by applying DBSCAN for the center points of the clusters, and the growth direction of each tree was estimated with principal component analysis (PCA). The points within accepted clusters were projected into the plane perpendicular to the growth direction for the final diameter estimation. The diameters were processed into a stem curve by automatically removing clear outliers, followed by fitting a smoothing spline to the remaining diameter estimates as a function of height. Finally, the DBH estimates were calculated by interpolating the stem curve at a height of 1.3 m above the ground. For a detailed description of the method, the reader is referred to hyyppa2020accurate; hyyppa2020under.
@@ 3.7.1 Autonomous flying performance
The accuracy of the VIO localization was evaluated by calculating the Absolute Trajectory Error (ATE), as proposed by sturm2012ate, with respect to the post-processed SfM reference trajectories. In this study, ATE was calculated only for the position estimate, and the error in the orientation estimate was not evaluated. ATE was calculated using the trajectory evaluation toolbox by rpgToolbox, which aligns the trajectory estimates with the reference trajectories with the method by Umeyma1991. After aligning the trajectories, the toolbox calculates the RMSE of the individual state errors to generate a single metric error value representing the accuracy of the entire trajectory. In the mathematical form, the error can be presented as
@@
In the first tests conducted in a simulator as described by (karjalainen2023), the VINS-Fusion estimate was found to be more accurate when the $t_{d}$ value was fixed to a constant, compared to when the online estimation method (onlineTd) was used. For this reason, during the test flights, the $t_{d}$ value was fixed to the one obtained from kalibr. However, the calibration estimate is never perfect, and the prototype did not have actual hardware synchronization between the camera and the autopilot IMU. For that reason, after the test flights, the VINS-Fusion estimation was also run with online estimated $t_{d}$ with the camera and IMU data recorded during the flight tests. The estimates were performed with the onboard computer of the drone, and the $t_{d}$ estimate obtained with kalibr was used as the initial value of the online estimation.